(1 point) Based on information from Diagnostic Tests with Nursing Implications, Springhouse Corporation, for the population of healthy female adults, the mean red blood cell count (RBC) is 4.8 (millions of cells per cubic millimeter of whole blood). A researcher thinks the RBC is underestimated, so she conducts a test of significance using Null Hypothesis: u s 4.8 vs. Alternative Hypothesis: ? > 4.8. If the true mean RBC is 4.75 and the null hypothesis is rejected, what has occurred? A. No error has occurred B. Type II Error has occurred C. An error in the z-test statistic occurred D. Type I Error has occurred E. Cant tell without knowing the degrees of freedom 8. 9. (1 point) statistic of 1.45. Would you have the same conclusion if a 05 as you would with an a 10? A. True B. False You conduct a right-tailed one sample proportion test (with n>30) and find a z-test 10. (1 point) Suppose we want to test the hypothesis H.. ? = 5 vs. ?#5. Based on a random 2.32, what will be the correct statement about sample of size 15 we obtain the test statistic the p-value for this test? A. p-value -0.02 B. 0.02 <p-value <0.05 C. p-value 0.99 D. 0.01<p-value <0.025 E. 0.05 p-value <0.01 LI.(I point) We want to make a comparison on the proportion of deaths due to lung cancer in the plant workers versus the proportion of deaths due to lung general population of chemical cancer for those working in a new chemical plant. If we draw a sample and find the of deaths due to lung cancer in the general population of chemical estimated at 12.8% (n 250). Suppose that of 100 deaths that plant workers nationwide is occurred in a new chemical plant, 22 were found to be due to lung cancer. We wish to determine whether there is a difference between the proportion of deaths from lung cancer in this plant and the proportion in the general population of chemical plant workers. The hypotheses for this test are Ho: p 0 vs. HA: P1-P2 * 0, Anda-005. What is the SE for the difference in proportions? A. -0.0013 B. 0.0017 C. 0.0465 D. 0.2037 E. 0.0004
Answer #1

(8)

This is the case of incorrect rejection of a null hypothesis which is actually true, so this is actually a type I error.

(9)

False, because when the value of significance level decreases, the chances of rejection of null hypothesis are also reduced.

(10)

The p-value for this test is:

p = 0.0359

So the correct option is:

(B) 0.02 < p < 0.05

(11)

First we calculate the weighted proportion as follows:

p = (p1*n1 + p2*n2)/(n1+n2) = (0.128*250 + 0.22*100)/(250+100) = 0.154

The standard error is calculated as:

SE = ( p*(1-p)/n1 + p*(1-p)/n2 )0.5 = ( 0.154*(1-0.154)/250 + 0.154*(1-0.154)/100 )0.5 = 0.0427

So the closest answer is 0.0465

Hope this helps !